OpenStudy (anonymous):

I have a question on my precal review for a test. The problem is: Show how(show work) CotӨ times SinӨ equals CosӨ (CotangentӨ * SineӨ= CosineӨ) If you have any idea how to work this or what this means please tell me. Any help would be greatly appreciated.

hartnn (hartnn):

do you know definition of cot , son and cos ratios ?

hartnn (hartnn):

**sin

OpenStudy (anonymous):

Cot equals 2squareroot5 over 5. Sin equals square root5 over 2. Cos equals 2 over 3. they are all in fraction form.

hartnn (hartnn):

where does these values come from ?? :O i assume you have to show that \(\cot \theta \times \sin \theta = \cos \theta \) start by drawing a right triangle,  i take angle B as theta